OpenStudy (anonymous):

A bus leaves the station at 19 mph. 6 hours later a car leaves the station traveling at 57 mph, and heads in the same direction. How far from the station will the two meet?

OpenStudy (anonymous):

first find the distance between them = 19*6 miles and since they are travelling in the same direction relative speed = 57-19 =38 mph time to meet is 19*6 /38 =3hr hence they will meet at 57*3=171 miles away from station

OpenStudy (anonymous):

in three hours after the car leaves they meet after 171 miles

OpenStudy (anonymous):